ok i need help on algebra i forgot what it is called but it is like 120% of 4/5
please please help me.

Found 2 solutions by checkley77, mathwiz687:
Answer by checkley77(12844)

120% of 4/5
1.2*4/5=4.8/5=48/50=24/25 answer.
or:
1.2*.8=.96 answer.

Answer by mathwiz687(1)

you go like this 120% is 120 over 100 of means multiply so it is 120 x 4 = 480
100 5 500